OpenStudy (anonymous):

simplify -3x^2 -2 = x^2 + 3x

OpenStudy (anonymous):

4x^2+3x+2=0 use the quadratic formula to find the roots...

OpenStudy (anonymous):

\[-4x ^{2}-3x-2=0\] you can re write it as \[4x ^{2}+3x+2=0\]